1. Uvita Associates acquired $8,000,000 par value, 11%, 4-year bonds on their date of issue, January 1 of the current year. The market yield at the time of issue is 10% for bonds of similar risk and maturity. Interest is paid semiannually on July 1 and January 1. Uvita uses the effective interest rate method to account for this investment. Uvita intends to keep the bonds available for sale. The fair value of the bonds at the end of the year of acquisition is $8,100,000. Minor rounding is ignored when grading your answers.

A. Determine the purchase price of the investment in bonds.

B. Prepare the journal entry to record the acquisition of the bond investment.

C. Prepare any needed year end journal entries pertaining to this investment. Be sure to identify any accounts that are OCI (Other Comprehensive Income) accounts. Include any closing entry for OCI accounts to AOCI.
Answer #1

Uvita Associates acquired $8,000,000 par value, 11%, 4-year bonds on their date of issue, January 1 of the current year.

The market yield at the time of issue is 10% for bonds of similar risk and maturity.

Interest is paid semiannually on July 1 and January 1.

Uvita uses the effective interest rate method to account for this investment.

Uvita intends to keep the bonds available for sale.

The fair value of the bonds at the end of the year of acquisition is $8,100,000.

A. Determine the purchase price of the investment in bonds.

purchase price = Present value of interest + present value of Par (maturity value )

Or

purchase price = Interest * PVIFA + Par * PVIF

Interest = 8000000 * 11% * 1 / 2 = 440000

PVIFA = Present value of interest factor annuity at 5% ( 10 / 2 ), 8 ( 4 * 2 ) = 6.46321

Par = 8000000

PVIF = present value interest factor at 5%, for 8 half year period = 0.676839

purchase price = 440000 * 6.46321 + 8000000 * 0.676839

purchase price = 2843812.4 + 5414712 = $8258524

B. Prepare the journal entry to record the acquisition of the bond investment.

Date

Accounts & Explanation

Debit

Credit

Jan 1

Investment Bond - AFS

$8000000

Premium on Bond investment

( 8258524 - 8000000)

$258524

Cash

$8258524

.

(AFS - Bond purchased at premium )

.

.

C. Prepare any needed year end journal entries pertaining to this investment. Be sure to identify any accounts that are OCI (Other Comprehensive Income) accounts. Include any closing entry for OCI accounts to AOCI.

Date

Cash received for interest

Interest Income

Amortization

Carrying amount

(8000000*5.5%)

(Carrying amount * 5%)

1/1

$8258524

6/30

440000

8258524*5%

=412926.2

27073.8

8231450.2

12/31

440000

8231450.2*5%

=411572.51

28427.49

8203022.71

Carrying value at end of the first year = 8203022.71

Fair value at end of the first year = 8100000

Fair value adjustment,

Unrealized loss on bond investment - OCI = 8203022.71 - 8100000 = 103022.71

Journal entry for fair value adjustment

Date

Accounts

Debit

Credit

Dec 31

Unrealized loss on bond investment - OCI

$103022.71

Investment Bond - AFS

$103022.71

.

(fair value adjustment )

.

.